# Probing Politico-Economic Bias in Multilingual Large Language Models: A Cultural Analysis of Low-Resource Pakistani Languages

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly shaping public discourse, yet their politico-economic biases remain underexamined in non-Western and low-resource multilingual contexts. This paper presents a systematic analysis of political bias in 13 state-of-the-art LLMs across five low-resource languages spoken in Pakistan—Urdu, Punjabi, Sindhi, Balochi, and Pashto. We propose a novel framework that integrates an adapted Political Compass Test (PCT) with a multi-level framing analysis. Our method combines quantitative assessment of political orientation across economic (left–right) and social (libertarian–authoritarian) axes with qualitative analysis of framing through content, style, and emphasis. We further contextualize this analysis by aligning prompts with 11 key socio-political themes relevant to Pakistani society. Our results reveal that LLMs predominantly align with liberal-left values, echoing Western training data influences, but exhibit notable shifts toward authoritarian framing in regional languages, suggesting strong cultural modulation effects. We also identify consistent model-specific "bias signatures" and language-conditioned variations in ideological expression. These findings show the urgent need for culturally grounded, multilingual bias auditing frameworks.
Probing Politico-Economic Bias in Multilingual Large Language Models:
A Cultural Analysis of Low-Resource Pakistani Languages
@@ 1 Introduction
Large Language Models (LLMs) have rapidly transformed the field of Natural Language Processing (NLP), demonstrating impressive capabilities across a wide array of linguistic and cultural contexts Blodgett et al. (2020). As these models become increasingly integrated into daily life, growing concerns have emerged about their potential to propagate social harms, including hallucinations, gender, and racial biases Zheng et al. (2023). Among these concerns, political bias stands out as particularly consequential, given its potential to misrepresent public discourse, reinforce dominant ideologies, and exacerbate societal polarization Demszky et al. (2019).
LLMs often reflect the socio-political assumptions embedded in their training data, frequently mirroring political biases aligned with high-resource, Western contexts Weidinger et al. (2021). While significant research has focused on political bias in models operating in English, Chinese, and other high-resource languages, relatively little attention has been paid to how such biases manifest in low-resource languages—especially in multilingual, culturally complex regions like South Asia Kumar et al. (2023). The intersection of linguistic diversity and socio-political dynamics in low-resource languages poses unique challenges for building fair and culturally attuned AI systems Suarez et al. (2019).
Pakistan presents a particularly compelling case for such investigation Harman (2018). With five major low-resource languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—spoken by over 80% of the population, Pakistan encompasses diverse political ideologies shaped by religion, regional history, and socio-economic disparities Sun et al. (2022). Sensitive topics such as blasphemy laws, minority rights, language policies, and provincial autonomy dominate political discourse, making it imperative to understand how LLMs engage with such issues across different linguistic and cultural dimensions.
Existing research on political bias in LLMs suffers from several important limitations Naveed et al. (2024). First, current studies largely focus on Western languages and political paradigms, leaving non-Western contexts critically underexplored Chen et al. (2020). Second, most work isolates stance detection from the broader question of how political orientations are linguistically framed in generated content Yu et al. (2023). Third, current evaluation frameworks rarely account for the cultural nuances and socio-political particularities that shape language use in South Asia Bordia and Bowman (2019). While some studies have examined multilingual LLMs, they often treat languages as independent variables rather than analyzing how political ideologies may transfer, transform, or intensify across culturally linked but linguistically distinct languages Bang et al. (2021).
This research addresses these gaps by introducing a comprehensive, culturally grounded framework for evaluating political and economic bias in multilingual LLMs, focusing specifically on five low-resource Pakistani languages. We propose a novel two-tier methodology that combines an adapted Political Compass Test (PCT) with a fine-grained framing analysis that examines both what is said (stance) and how it is said (style), as shown in Figure 1. Our approach captures ideological bias not only in the LLMs’ political positions but also in their narrative framing, across languages that reflect diverse cultural and political realities. This study makes the following key contributions:
@@
  • We conduct the first large-scale evaluation of political bias in LLMs across five low-resource Pakistani languages.
  • We adapt and manually translate the PCT into these languages and align it with 11 culturally salient topics such as LGBTQ+ rights, blasphemy laws, and language policy, grounded in the Pakistani socio-political context.
  • We introduce a three-pronged framing analysis that integrates Boydstun’s frame dimensions, NER, and lexical polarity assessment to dissect both content and stylistic framing biases.
@@ 2 Related Work
Political Bias in Language Models: The political orientation of LLMs has become a core issue in AI ethics. Early work by Liu et al. (2021) showed that word embeddings reflect human-like ideological bias. Later studies, such as Motoki et al. (2024) and Ceron et al. (2024), demonstrated that models like GPT-3/4 tend to lean liberal on social matters and can convincingly emulate partisan ideologies. However, these findings are largely based on U.S.-centric frameworks, limiting their cross-cultural relevance.
Evaluation Tools: Political Compass and Policy Probes: The Political Compass Test (PCT) has been widely used for evaluating AI political alignment Hartmann et al. (2023), offering a structured assessment across ideological axes. Others like Bang et al. (2021) introduced alternative methods based on policy preference tasks. Yet, these frameworks are developed for English-speaking contexts, often assuming direct translatability across languages and cultures—an assumption that fails in multilingual and low-resource settings.
Bias in Multilingual and Low-Resource Contexts: Research has shown that traditional bias detection tools underperform in non-English environments due to linguistic and cultural misalignment Barkhordar et al. (2024). South Asian languages such as Punjabi, Sindhi, Pashto, and Balochi are especially underserved in LLM training data. In other study Harman (2018) further found that sentiment models trained on English data degrade when applied to translated content, underscoring the need for culturally-aware evaluation strategies. Recent work also cautions against treating multilingual bias as language-independent Thapa et al. (2024).
Framing and Discourse-Level Analysis: Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing. The taxonomy by Boydstun et al. (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations. Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure—areas that remain largely unexplored in multilingual NLP.
@@
Despite recent progress, several critical limitations remain in the study of political bias in language models. Much of the existing literature is Western-centric, focusing predominantly on high-resource languages like English and Spanish, with limited attention to multilingual or non-Western contexts. Moreover, while stance detection is a common approach, it often overlooks how political bias is conveyed through rhetorical framing. Tools like the PCT, though widely used, have not been culturally adapted for regions like South Asia, limiting their interpretability and relevance. In response to these gaps, our study makes several key contributions. We culturally adapt the PCT for five low-resource Pakistani languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—ensuring linguistic accuracy and ideological relevance. We propose a novel three-layered framing analysis that integrates policy frame taxonomy, named entity recognition (NER), and lexical polarity scoring to capture both content and style. Our evaluation spans 11 culturally salient political topics central to Pakistani discourse, and we benchmark 13 leading language models, both open-source and commercial. To our knowledge, this is the first comprehensive investigation of political bias in LLMs conducted within a multilingual, non-Western framework.
@@ 3.1 Political Leaning of pretrained LLM
Overview and Motivation: To evaluate political bias in large language models (LLMs), we adopt a two-step evaluation approach grounded in political spectrum theories, following the framework proposed by Feng et al. (2023). This framework characterizes political opinion using a two-axis model: social values, ranging from liberal to conservative ($S_{\mathrm{soc}}$), and economic values, ranging from left to right ($S_{\mathrm{eco}}$). This dual-dimensional representation offers a nuanced view of political ideologies, surpassing the simplistic left-right dichotomy and enabling deeper analysis of how political bias manifests in LLMs Liu et al. (2024).
@@
Political Compass Test (PCT): The PCT [^1] https://www.politicalcompass.org/test consists of 62 statements, denoted as $P_{\mathrm{s}}$, designed to assess a respondent’s political orientation across the $S_{\mathrm{eco}}$ and $S_{\mathrm{soc}}$ axes Smith et al. (2022). Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree. The final scores are calculated based on weighted responses and mapped to a coordinate in the range $[-10,10]$ for each axis, producing a two-dimensional representation of political ideology.
Translation and Human Annotation: To adapt the PCT for low-resource languages in Pakistan, we employed native speakers with expertise in political discourse from five different cities. Each statement was translated into five Pakistani languages, detailed Table in Appendix 18: Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.
@@
Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models. The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o [^2] https://openai.com/api, Claude [^3] https://www.anthropic.com, and Gemini 1.5 Pro [^4] https://ai.google.dev/gemini-api/docs. The open-source models include: Mistral-7B-Instruct [^5] https://docs.mistral.ai/api/, DeepSeek [^6] https://lambda.ai/, RoBERTa-large [^7] https://huggingface.co/FacebookAI/xlm-roberta-large, RoBERTa-base11, and BERT-large 12. For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 15.
 [^8] https://huggingface.co/FacebookAI/xlm-roberta-base
Response Generation and Parameters: To ensure consistency, we fixed generation parameters across all models: $\text{Temperature}=0.5$, Top-$p$ = 1. All responses were generated in a zero-shot multilingual setting. For stance detection, we used the multilingual mDeBERTa-v3-base model, which supports over 100 languages, including all five under study.
@@
Stance Scoring and Mapping: To quantify the model’s political stance on each PCT statement, we employ a two-stage computational approach. First, we calculate a scalar stance score from classifier confidence values. Second, we discretize this continuous score into ordinal categories that align with the PCT response format. This allows us to convert raw classifier outputs into interpretable symbolic judgments.
Let $A_{s},A,D,D_{s}\in[0,1]$ denote the predicted confidence scores corresponding to the classes Strongly Agree, Agree, Disagree, and Strongly Disagree, respectively.
@@
1. Stance Score Computation: We define a scoring function $f:[0,1]^{4}\rightarrow[-10,10]$ to assign a value based on the dominant confidence class:
@@
The result $S=f(A_{s},A,D,D_{s})\in[-10,10]$ serves as a continuous stance score, indicating both the direction and strength of agreement.
2. Stance Discretization: To facilitate comparative analysis across models and statements, we define a discretization function $g:[-10,10]\rightarrow\{0,1,2,3\}$ that maps the continuous score to categorical labels using a symmetric threshold parameter $\tau>0$:
$g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$
This results in an ordinal stance label interpreted as:
  • 3 – Strongly Agree
  • 2 – Agree
  • 1 – Disagree
  • 0 – Strongly Disagree
By mapping soft classifier outputs to these well-defined categories, we ensure that downstream aggregation and political leaning visualization remain interpretable and robust. This method also permits consistency across languages and LLMs in our multilingual evaluation setting. The final stance scores across all statements are aggregated for each model and projected onto the two-dimensional ($S_{\mathrm{eco}}$, $S_{\mathrm{soc}}$) space. This facilitates a structured evaluation of political alignment and model behavior across both ideological dimensions.
@@ 3.2 Ideological Framing Analysis
Each topic was translated into five low-resource Pakistani languages to enable multilingual evaluation.
For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$. Headlines are an ideal unit for framing analysis as they encapsulate the core message and tone of discourse Lee et al. (2022); Sheng et al. (2021); Baly et al. (2020). We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al. (2021) (see Appendix  15 for prompt details).
Frame Dimension Classification: To investigate the ideological narratives, we classify generated headlines using Boydstun’s 15 cross-cutting frame dimensions Boydstun et al. (2014), which capture salient political discourse themes such as “Economics,” “Morality,” “Health and Safety,” and “Cultural Identity.” These topic-agnostic frames facilitate comparative analysis across models and topics Hamborg (2020), for details, see Appendix D. Frame classification was performed using GPT-3.5-turbo with bilingual prompts, computing frame ratios to identify dominant framing strategies employed by each model Ziems and Yang (2021).
Entity-Based Framing Analysis: Beyond content frames, we analyze framing at the entity level by extracting and quantifying frequently mentioned named entities (e.g., political figures, countries, organizations) from generated headlines. Frequent entity mentions serve as a subtle framing device, indicating which actors or groups models emphasize Devlin et al. (2019); Zheng et al. (2023). NER was conducted using a pretrained bert-base-multilingual-cased model  [^11] https://huggingface.co/google-bert/bert-base-multilingual-cased.
Sentiment Polarity Towards Entities: To assess stylistic and attitudinal bias, we analyze the sentiment polarity expressed towards entities within headlines. Using a fine-tuned XLM-RoBERTa model for Urdu sentiment analysis Fan et al. (2019), we categorize sentiment as positive, negative, or neutral Saez-Trumper et al. (2013). Aggregating sentiment across samples yields a profile of the model’s overall stance tone towards each topic, revealing subtle stylistic bias Spliethöver et al. (2022); Roy and Goldwasser (2020).

### v2

# Framing Political Bias in Multilingual LLMs Across Pakistani Languages

@@ Abstract
Large Language Models (LLMs) increasingly shape public discourse, yet most evaluations of political and economic bias have focused on high-resource, Western languages and contexts. This leaves critical blind spots in low-resource, multilingual regions such as Pakistan, where linguistic identity is closely tied to political, religious, and regional ideologies. We present a systematic evaluation of political bias in 13 state-of-the-art LLMs across five Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi. Our framework integrates a culturally adapted Political Compass Test (PCT) with multi-level framing analysis, capturing both ideological stance (economic/social axes) and stylistic framing (content, tone, emphasis). Prompts are aligned with 11 socio-political themes specific to the Pakistani context. Results show that while LLMs predominantly reflect liberal-left orientations consistent with Western training data, they exhibit more authoritarian framing in regional languages, highlighting language-conditioned ideological modulation. We also identify consistent model-specific bias patterns across languages. These findings show the need for culturally grounded, multilingual bias auditing frameworks in global NLP.
[…]
202F
Framing Political Bias in Multilingual LLMs Across Pakistani Languages
@@ 1 Introduction
Large Language Models (LLMs) have achieved strong performance across a range of NLP tasks and languages Blodgett et al. (2020). However, increasing evidence shows that these models encode social and ideological biases, including hallucinations, stereotypes, and political partisanship Zheng et al. (2023). Political bias is particularly consequential, as it can influence public discourse, reproduce dominant ideologies, and marginalise minority perspectives Demszky et al. (2019).
Bias in language models is not a theoretical concern—it shapes real-world outputs such as news headlines, reinforcing dominant ideologies while marginalizing dissent. This can distort public discourse, erode democratic values, and undermine trust in AI systems Barkhordar et al. (2024). While English and other high-resource languages have received some scrutiny Weidinger et al. (2021), low-resource languages like Urdu, Punjabi, Sindhi, Pashto, and Balochi remain severely underexplored Kumar et al. (2023). In Pakistan, where political identity is closely tied to language, this oversight risks amplifying bias across culturally sensitive issues. Addressing this gap is vital to building fair, inclusive AI systems for linguistically diverse and politically complex societies.
Figure 1: An illustration of political bias in multilingual LLMs using a real Urdu-language response to PCT Statement 24. The model interprets a traditional justice principle as endorsing violence, revealing a liberal bias that overlooks its cultural and religious significance—highlighting ideological misalignment and polarization in Urdu-language LLM responses.
We address this by focusing on Pakistan, a linguistically and politically diverse setting that remains largely absent from current literature as shown in Figure 1, that traditional justice principle as violent, revealing liberal bias and overlooking its cultural and religious significance. Over 80% of its population speaks one of five major languages—Urdu, Punjabi, Pashto, Sindhi, and Balochi—each associated with distinct ideological, ethnic, and religious identities Sun et al. (2022). Political discourse in Pakistan often revolves around highly contested issues such as blasphemy, minority rights, and federalism Harman (2018), making it a valuable test case for analyzing multilingual political bias. Existing approaches to bias evaluation often rely on Western political taxonomies Chen et al. (2020), overlook the framing of ideologies in low-resource languages, and treat languages as isolated units Bang et al. (2021). Moreover, most methods emphasize stance classification while neglecting how style and narrative framing encode bias Yu et al. (2023). Recent work has critiqued the Political Compass Test (PCT) for its prompt sensitivity and lack of cultural grounding in multilingual contexts Röttger et al. (2024), calling for more context-aware evaluations. Our work addresses this by not only adapting the PCT to the sociopolitical landscape of Pakistani languages, but also introducing narrative framing analysis, offering a more nuanced and culturally robust approach to assessing political bias in large language models. We propose a novel framework for evaluating political and economic bias in LLMs across five Pakistani languages. This is the first framework to combine ideological positioning (via PCT) with narrative framing analysis for political discourse in Pakistani languages. This study makes the following contributions:
@@
  • We conduct the first large-scale political bias evaluation in five Pakistani languages.
  • We adapt and translate the PCT to cover 11 culturally salient topics grounded in Pakistani discourse.
  • We propose a three-part framing analysis using Boydstun’s taxonomy, named entity recognition, and lexical polarity.
@@ 2 Related Work
2.1 Political Bias in Language Models:
The political orientation of LLMs has emerged as a core concern in AI ethics. Studies show models like GPT-3/4 reflect liberal social leanings and partisan patterns Liu et al. (2021); Motoki et al. (2024); Ceron et al. (2024), but these insights remain western-centric. Tools like the Political Compass Test (PCT) Hartmann et al. (2023) and policy probes Bang et al. (2021) assume linguistic translatability, which fails in low-resource contexts. Bias detection tools often underperform in non-English settings due to cultural misalignment Barkhordar et al. (2024). Pakistani languages—Punjabi, Sindhi, Pashto, and Balochi—are critically underserved, requiring culturally grounded evaluation strategies Harman (2018); Thapa et al. (2024).
2.2 Framing and Discourse-Level Analysis:
Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing. The taxonomy by Boydstun et al. (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations. Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure—areas that remain largely unexplored in multilingual NLP.
@@ 2.2 Framing and Discourse-Level Analysis:
2.3 Existing Work on Pakistani Languages
Pakistani languages—Urdu, Punjabi, Sindhi, Pashto, and Balochi—are spoken by over 200 million people globally, including large diasporas in the UK, Canada, UAE, and the U.S. Mostefa et al. (2012); Hussain (2004). Despite this, they remain critically under-represented in NLP. Recent work has addressed Urdu QA Arif et al. (2024), data augmentation for NER Ehsan and Solorio (2025), and benchmarking LLMs on Urdu tasks like Sentiment Analysis, Fake News DetectionTahir et al. (2025). However, political bias and framing remain unexplored. Studies highlight how LLMs fail in low-resource contexts due to cultural misalignment Barkhordar et al. (2024); Thapa et al. (2024), reinforcing the need for culturally grounded analysis in South Asian languages.
Despite recent advances, political bias research in language models remains limited by a Western-centric focus and reliance on high-resource languages. Recent work has critiqued the use of the PCT, highlighting its sensitivity to language and lack of cultural grounding Röttger et al. (2024). Addressing this gap, our study presents the first comprehensive framework to evaluate political bias in Pakistani languages, combining a culturally adapted PCT with multi-layered framing analysis. Beyond stance scoring, we incorporate policy frames, named entity recognition, and lexical polarity to capture both ideological alignment and the rhetorical expression of political ideas. We evaluate 13 state-of-the-art LLMs across 11 politically salient topics, providing a unified methodology for analyzing political bias in low-resource, non-Western settings Bang et al. (2024).
@@ 3.1 Political Leaning of pretrained LLM
Overview and Motivation: To evaluate political bias in large language models (LLMs), we adopt an evaluation approach grounded in political spectrum theories, following the framework proposed by Feng et al. (2023). This framework characterizes political opinion using a two-axis model: social values, ranging from liberal to conservative ($S_{\mathrm{soc}}$), and economic values, ranging from left to right ($S_{\mathrm{eco}}$). This representation offers a nuanced view of political ideologies, surpassing the simplistic left-right dichotomy and enabling deeper analysis of how political bias manifests in LLMs Liu et al. (2024).
@@
Political Compass Test (PCT): The PCT [^1] https://www.politicalcompass.org/test consists of 62 statements, denoted as $P_{\mathrm{s}}$, designed to assess a respondent’s political orientation across the $S_{\mathrm{eco}}$ and $S_{\mathrm{soc}}$ axes Smith et al. (2022). Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree. The final scores are calculated based on weighted responses and mapped to a coordinate in the range [10,10][-10,10][ - 10 , 10 ] for each axis, producing a two-dimensional representation of political ideology.
@@
Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models. The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o, Claude, and Gemini 1.5 Pro. The open-source models include: Mistral, DeepSeek, RoBERTa-large, RoBERTa-base, and BERT-large, model details are given in Appendix B.3. For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 23.
@@
Stance Scoring and Mapping: To quantify the model’s political stance on each PCT statement, we employ a two-stage computational approach. First, we calculate a scalar stance score from classifier confidence values, for more details see Appendix B.4. Second, we discretize this continuous score into ordinal categories that align with the PCT response format. We discretize stance by selecting the class with the highest softmax confidence score among four candidate labels. The continuous stance score is then calculated by multiplying the confidence value with a fixed weight: $\pm$10 for Strongly Agree/Disagree and $\pm$5 for Agree/Disagree. This approach reflects both the categorical stance and the model’s certainty. This allows us to convert raw classifier outputs into interpretable symbolic judgements. Let $A_{s},A,D,D_{s}\in[0,1]$ denote the predicted confidence scores corresponding to the classes Strongly Agree, Agree, Disagree, and Strongly Disagree, respectively.
@@
1. Stance Score Computation: We define a scoring function $f:[0,1]^{4}\rightarrow[-10,10]$ to assign a value based on the dominant class:
@@
The result $S=f(A_{s},A,D,D_{s})\in[-10,10]$ serves as a continuous stance score, indicating both the direction and strength of agreement. 2. Stance Discretization: To facilitate comparative analysis across models and statements, we define a discretization function $g:[-10,10]\rightarrow\{0,1,2,3\}$ that maps the continuous score to categorical labels using a symmetric threshold parameter $\tau>0$:
$g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$ (1)
[…]
@@ 3.2 Ideological Framing Analysis
Each topic was translated into five Pakistani languages to enable multilingual evaluation. For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$. Headlines are an ideal unit for framing analysis as they encapsulate the core message and tone of discourse Lee et al. (2022); Sheng et al. (2021); Baly et al. (2020). We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al. (2021) (see Appendix 23 for prompt details and Appendix B.2for prompting strategy and reproducibility details).
Frame Dimension Classification: To examine ideological narratives in model-generated content, we classify headlines using Boydstun’s 15 cross-cutting frame dimensions (Boydstun et al., 2014), which encompass salient themes such as Economics, Morality, Health and Safety, and Cultural Identity. These topic-independent frames enable consistent comparative analysis across models and topics (Hamborg, 2020). We employ GPT-3.5-turbo with bilingual prompts to classify each headline into one or more frames, enhancing contextual understanding in Pakistani languages. For each topic–stance pair $(t,s)$, we compute the frame ratio for frame $f_{i}$ as:
$\text{FrameRatio}_{t,s}(f_{i})=\frac{c_{t,s}(f_{i})}{N_{t,s}}$ (2)
[…]
Entity-Based Framing Analysis: To explore how models frame specific actors or institutions, we extract named entities—such as political figures, countries, and organizations—from Urdu headlines using a multilingual NER model. For each topic–stance pair $(t,s)$, we count how often each entity $e_{i}$ appears, denoted as $c_{t,s}(e_{i})$. We then calculate its relative Prominence P using:
$\text{P}_{t,s}(e_{i})=\frac{c_{t,s}(e_{i})}{\sum_{j=1}^{n}c_{t,s}(e_{j})}$ (3)
This score reflects which entities are most emphasized in model outputs, offering insights into how narratives center around particular individuals or groups. Frequent entity mentions serve as a subtle framing device, indicating which actors or groups models emphasize Devlin et al. (2019); Zheng et al. (2023). NER was conducted using a pretrained bert-base-multilingual-cased model  [^4] https://huggingface.co/google-bert/bert-base-multilingual-cased.
[…]
$\Pr^{(k)}_{t,s}(e_{i})=\frac{S_{e_{i}}^{(k)}}{\sum_{k^{\prime}}S_{e_{i}}^{(k^{\prime})}}$ (4)
We define the dominant sentiment polarity for entity $e_{i}$ as:
$\text{SentimentBias}_{t,s}(e_{i})=\arg\max_{k}\left(\Pr^{(k)}_{t,s}(e_{i})\right)$ (5)
These scores expose entity-level sentiment bias patterns across topics and stances (Spliethöver et al., 2022; Roy and Goldwasser, 2020).